36 boys in a class weigh 42 kg on an average . if their teacher is included, the average increase by 1.5 kg, find the teachers weigh
Answer by ssenin(1)

x= all boys's weight
y= all boys and teacher weight
x/36 = 42
x = 1,512
y/37 = 43.5
y = 1,609.5
y-x = teacher's weight = 1,609.5-1,512 = 97.5